• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name D. McNulty

    Log entry time 19:05:34 on January31,2008

    Entry number 228410

    keyword=strange multiple peaks in thin detector with thin trigger

    After establishing the thin detector trigger (T3), we took a few runs: 
    3730 t3 trig, 260Hz T3 rate (very similar to t4 rate), 15mV T3 thresh
    3733 same as above with 16ns gate delay
    3735 same as above with 40mV threshold on T3
    
    The thin spectra from these runs all look like the left plot (with strange 
    multiple peaks)--we see these peaks in both thin and stack.  The two plots 
    below compare the thin spectra (triggered on the stack) after (left) and before 
    (right) the thin trigger was established.  The left plot is from run 3736, the 
    right plot is from run 3720.  We expected the left plot to look identical to 
    the right plot (same trigger conditions,etc), but this is not the case.  We 
    suspect the noisy fan-out problem could be causing this but are not sure.  We 
    made an access to undo the thin trigger electronics and return to previous 
    state.  
    	Alex will examine the scope trace of thin detector to determine if the 
    electronic's noise could be the culprit.  We will then switch over to 
    integration mode (during the same access) if we are confident we can reproduce 
    the original (right plot) spectra.
    
    Having said this, it seems we are abandoning our attempt to assess any 
    big-pulses in the thin detector.
    


    FIGURE 1

    FIGURE 2